OpenStudy (gabylovesyou):

4(x + 1) = 4x + 4 Solve for x! A. No solution B. 0 C. Any real number D. 1

jimthompson5910 (jim_thompson5910):

what do you get when you solve for x

OpenStudy (anonymous):

4x+4 = 4x+4 4x-4x =4-4 0=0 D. 1

jimthompson5910 (jim_thompson5910):

it's not D

OpenStudy (gabylovesyou):

its not b or d

jimthompson5910 (jim_thompson5910):

what is 4(x+1)

OpenStudy (gabylovesyou):

i say its C

jimthompson5910 (jim_thompson5910):

what does it distribute to

jimthompson5910 (jim_thompson5910):

C is correct

OpenStudy (anonymous):

@jim_thompson5910 what do you suppose it is?

OpenStudy (gabylovesyou):

yaaaaay!!!!!!!! lol

jimthompson5910 (jim_thompson5910):

any real number works

jimthompson5910 (jim_thompson5910):

plug in any number you want for x, and it will work

jimthompson5910 (jim_thompson5910):

this is because 4(x+1) is the same as 4x+4

OpenStudy (anonymous):

@jim_thompson5910 true